# Evaluation, Analysis and Mitigation in Multimodal Large Language Models

### v1

@@ 4 Analysis of Relation Hallucination
Table 4: The average probability of all hallucination cases. We test LLaVA and MiniGPT4-v2 on Reefknot and two other representative relation hallucination benchmarks.
@@
Figure 7: The respective ratios between hallucination and non-hallucination with different entropy values. We conducted experiments on POPE, MMRel and our Reefknot.
To investigate the mechanism behind hallucination generation, we conducted an in-depth analysis of the fluctuations in probability values across model layers. MLLMs consist of a vision encoder, a projection layer, and a strong LLM decoder, which is stacked by a set of $\mathcal{N}$ transformer layers, and an MLP layer $\phi(\cdot)$ for predicting the next-word probability of distribution auto-regressively. Given an image represented by vision encoder as $\mathcal{V}_{t}$ = $\{v_{1},v_{2},\dots,v_{t}\}$ and a text prompt of tokens $\mathcal{P}_{t}$ = $\{p_{1},p_{2},\dots,p_{t}\}$, they are processed as a sequence $\mathcal{H}_{0}$ = $\psi(\mathcal{V}_{t},\mathcal{P}_{t})$ through projection and concatenation function $\psi(\cdot)$. Thus $\mathcal{H}_{0}=\{h_{1}^{(0)},\dots,h_{t-1}^{(0)}\}$, in which $h_{i}^{(k)}$ means the hidden states of $i_{th}$ token in $k_{th}$ language decoder layer.
Then $\mathcal{H}_{0}$ would be processed by each of the transformer layers in the language decoder successively. We denote the output of the $j$-th layer as $\mathcal{H}_{j}$. In the normal forward process, $\mathcal{H}_{j}$ will be calculated by $\mathcal{N}$ times, then it will be passed through language model head layer $\phi(\cdot)$ to predict the probability of the next token $r_{t}$ over the vocabulary set $\mathcal{X}$. In our set, we manually pass every $\mathcal{H}_{n}$ to explore the mechanism of hallucination with probability distributions and the next token to generate. For every layer, we can obtain the probability distribution $\mathbb{P}$ and the next word prediction $r_{t}^{(j)}$.
@@
Using Equation 4, we visualized these changes during the forward propagation process, as illustrated in Figure 8. To avoid variability from individual instances, we reported the average values of all data involving hallucinations occurring in Reefknot. Furthermore, to ensure a fair comparison, we employed both the 32-layer MiniGPT4-v2-7B (Chen et al. 2023) and the 40-layer LLaVA-13B (Liu et al. 2023a).
Figure 8: The average probability among layers when hallucinations occur. Sharp changes in probability happen in deep layers. Note we report MiniGPTv2-7B with 32 layers and LLaVA-13B with 40 layers to prove universality.
@@ 5 Detect-Then-Calibrate Mitigation Method
The evaluation results indicate that MLLMs commonly suffer from severe relation hallucinations. As analyzed in Section 4, we found that these relation hallucinations primarily stem from the model’s lack of confidence. The lack of confidence in the model’s responses results in a relatively high entropy value. Therefore, we can detect the occurrence of relation hallucinations by monitoring changes in entropy.
Initially, we set a specific entropy threshold $\gamma$ to detect potential hallucinations in the model’s output. If the entropy of the model’s response exceeds $\gamma$, which suggests a significant lack of confidence, we infer a high probability that the model has generated hallucinations. In cases where the model is identified to hallucinate, we will utilize the hidden states from intermediate layers to calibrate the final outputs layers, which is inspired by DoLa (Chuang et al. 2023) approach. Note, unlike traditional contrastive decoding strategies (Li et al. 2023a; Leng et al. 2023; Huang et al. 2023), we do not calibrate all cases. Instead, we focus on mitigating potential hallucination cases to avoid altering non-hallucinatory cases into hallucination ones. Formula 5 show our calibration process.
@@
Note that $\lambda$ is the hyperparameter to control the degree of the intermediate layer. The overall algorithmic flow can be seen from the pseudo-code below:
@@
As illustrated in Table 5, we conducted experiments using LLaVA-13B. To demonstrate the robustness of our results, we employed two additional relation hallucination datasets, MMRel (Nie et al. 2024) and R-bench (Wu et al. 2024). During experiments, we set $\lambda$ = 2 ,$\alpha$ = 0.1, $\gamma=0.9$. For a fair comparison, in addition to reporting the baseline model, we also report some training-free methods such as VCD (Leng et al. 2023), DoLa (Chuang et al. 2023), and OPERA (Huang et al. 2023). Our approach achieved improvements across all three relation hallucination datasets. Specifically, on the MMRel dataset, our model achieved a 19.7% improvement compared to the baseline setting.
@@
+ VCD (Leng et al. 2023) 38.32 41.96 22.05
+ DoLa (Chuang et al. 2023) 36.96 39.68 23.52
+ OPERA (Huang et al. 2023) 35.73 39.22 26.73
@@
Table 5: The hallucination rates of LLaVA-13B and its variants with hallucination mitigation techniques across Reefknot and two other relation hallucination benchmarks.
@@ 6 Case study
To intuitively demonstrate the generation process of hallucination, we visualized a real case from Reefknot in Figure 9. In the image, when we ask “Is the boy eating pizza in the photo?”, MLLMs are unable to provide precise answers. To investigate this, we analyze the probabilities associated with the top five predicted tokens at each layer of the language model. As illustrated in Figure 9, the model typically converges on the correct answer as it progresses from shallow to deep layers. Notably, hallucinations didn’t occur until the final layer.
Additionally, we can observe that as the number of layers increases, the model increasingly focuses on narrowing down the range of answer choices, with the probability for both ‘yes’ and ‘no’ rising.
However, as depicted in the changes of colors, when it reaches the final Transformer block (layer 38-40th), the model’s choices suddenly become uncertain and entropy rises, accompanied by the emergence of hallucinations. So we used logits of the intermediate layers, which are less likely to hallucinate, to calibrate the final logits to correct the answer. Following our calibration utilizing Equation 5, we not only dispelled any misconceptions but also significantly bolstered the confidence of our responses.
@@ 7 Related Work
Large Language Models (LLMs) are powerful tools, and exploring their interpretability (Han et al. 2024; Huo et al. 2024b), operational mechanisms, and methods to ensure trustworthy responses(Hu et al. 2023) has become a crucial area of research. In the realm of MLLMs, hallucinations are typically categorized into three distinct types: object, attribute, and relation, as outlined in prior studies (Yin et al. 2024; Liu et al. 2024). At the object and attribute levels, a considerable number of representative benchmarks such as POPE (Li et al. 2023c) and HaELM (Wang et al. 2023b) have been introduced by researchers. Many benchmarks are accompanied by a diverse array of evaluation criteria, with generative criteria including CHAIR (Li et al. 2023b), THRONE (Kaul et al. 2024) and various discriminative criteria being particularly notable.
The existing relation hallucinations benchmarks (Wu et al. 2024; Nie et al. 2024) focus only on discriminative criteria. The gap of current benchmarks is to identify relationship pairs that accurately reflect the dynamic interactions. Thus, we propose the Reefknot benchmark, incorporating both generative and discriminative criteria to comprehensively evaluate the relation hallucination.
@@
The concept of scene graphs was first introduced by Zhu et al. (2022), offering a novel method to describe and retrieve complex scenes within images. Building on this foundation, Krishna et al. (2016) further advanced the field by introducing the Visual Genome (VG) dataset in 2017. In recent years, there has been significant progress in leveraging this triplet-based structure for various applications. For instance, Yao et al. (2018) utilized scene graphs to generate more detailed image descriptions.
@@
Confidence estimation (Zou et al. 2023; Chen et al. 2024a) and calibration are essential for enhancing the reliability of LLMs such as GPT-3 (Brown et al. 2020). To assess the confidence associated with outputs from LLMs, Kuhn, Gal, and Farquhar (2023) have developed a method called semantic entropy that utilizes linguistic invariances reflecting shared meanings. However, this method relies on accessing token-level probabilities, which are often unavailable through current black-box APIs. Kadavath et al. (2022) have designed prompts that encourage the models to self-assess their responses and to explicitly calculate the probability that an answer is true; while Lin, Hilton, and Evans (2022) have prompted LLMs to provide both an answer and an accompanying confidence level. Additionally, Manakul, Liusie, and Gales (2023) introduced a sampling-based method to identify hallucinated facts in model outputs. In a similar vein, Kadavath et al. (2022) have designed prompts that encourage the models to self-assess their responses and to explicitly calculate the probability that an answer is true, further fine-tuning the model to improve the accuracy of these probabilities. Concurrently, Lin, Hilton, and Evans (2022) have prompted LLMs to provide both an answer and an accompanying confidence level. Manakul, Liusie, and Gales (2023) introduced a sampling-based method to identify hallucinated facts in model outputs.
8 Conclusion and Future Work
In conclusion, we propose a comprehensive benchmark called Reefknot to evaluate and mitigate relation hallucinations in MLLMs. We construct the dataset with over 20k data through a scene graph-based construction pipeline, covering two discriminative tasks (Y/N and MCQ) and one generative task (VQA). Our in-depth evaluation highlights a substantial performance gap on relation hallucination in existing MLLMs, emphasizing the need for more sophisticated reasoning capabilities. Subsequently, we discover that relation hallucinations tend to occur when MLLMs respond with low confidence. Therefore, we propose a Detect-then-Calibrate method to mitigate the relation hallucination via entropy threshold, with an average reduction of 9.75% in the hallucination rate across Reefknot and two other representative relation hallucination datasets. In general, we anticipate that our proposed Reefknot benchmark will pave the way for future advancements in trustworthy multimodal intelligence.
@@ 8 Conclusion and Future Work
Appendix
Detailed Case Study
@@
Manual Checking Rules
@@
More Error Analysis
@@
Visualization of Relation Word
@@
VQA Criterion
Here are the evaluation criteria for VQA questions. We use the Deberta model to determine whether the models entail each other. Only when the label and response contain each other will it be judged as a correct reply. We show our key function code in Figure 15. Besides, we show a VQA case in Figure 17.
More Cases
@@
3 #label is groundtruth;responese is the respone of LLM;

### v2

@@ 4 Analysis of Relation Hallucination
Table 4: The average probability of all hallucination cases. We evaluate LLaVA and MiniGPT4-v2 on the Reefknot dataset, along with two other representative relation hallucination benchmarks.
@@
Figure 6: The respective ratios between hallucination and non-hallucination with different entropy values.
To investigate the mechanism behind hallucination generation, we conducted an in-depth analysis of the fluctuations in probability values across model layers. MLLMs consist of a vision encoder, a projection layer, and a strong LLM decoder, which is stacked by a set of $\mathcal{N}$ transformer layers, and an MLP layer $\phi(\cdot)$ for predicting the next-word probability of distribution auto-regressively. Given an image represented by vision encoder as $\mathcal{V}_{t}$ = $\{v_{1},v_{2},\dots,v_{t}\}$ and a text prompt of tokens $\mathcal{P}_{t}$ = $\{p_{1},p_{2},\dots,p_{t}\}$, they are processed as a sequence $\mathcal{H}_{0}$ = $\psi(\mathcal{V}_{t},\mathcal{P}_{t})$ through projection and concatenation function $\psi(\cdot)$. Thus $\mathcal{H}_{0}=\{h_{1}^{(0)},\dots,h_{t-1}^{(0)}\}$, in which $h_{i}^{(k)}$ means the hidden states of $i_{th}$ token in $k_{th}$ language decoder layer. Then $\mathcal{H}_{0}$ would be processed by each of the transformer layers in the language decoder successively. We denote the output of the $j$-th layer as $\mathcal{H}_{j}$. In the normal forward process, $\mathcal{H}_{j}$ will be calculated by $\mathcal{N}$ times, then it will be passed through language model head layer $\phi(\cdot)$ to predict the probability of the next token $r_{t}$ over the vocabulary set $\mathcal{X}$. In our set, we manually pass every $\mathcal{H}_{n}$ to explore the mechanism of hallucination with probability distributions and the next token to generate. For every layer, we can obtain the probability distribution $\mathbb{P}$ and the next word prediction $r_{t}^{(j)}$.
@@
Using Equation 4, we visualized these changes during the forward propagation process, as illustrated in Figure 7. To avoid variability from individual instances, we reported the average values of all data involving hallucinations occurring in Reefknot. To ensure a fair comparison, we utilized the 32-layer MiniGPT4-v2-7B model Chen et al. (2023) and the 40-layer LLaVA-13B model Liu et al. (2023a).
Figure 7: The average probability across layers during hallucination occurrences. Significant probability changes are observed in deeper layers. Results are reported for MiniGPTv2-7B (32 layers) and LLaVA-13B (40 layers) to demonstrate universality.
@@ 5 Detect-Then-Calibrate Mitigation Method
The evaluation results indicate that MLLMs commonly suffer from severe relation hallucinations. As analyzed in Section 4, we found that these relation hallucinations primarily stem from the model’s lack of confidence. The lack of confidence in the model’s responses results in a relatively high entropy value. Therefore, we can detect the occurrence of relation hallucinations by monitoring changes in entropy. Initially, we set a specific entropy threshold $\gamma$ to detect potential hallucinations in the model’s output. If the entropy of the model’s response exceeds $\gamma$, which suggests a significant lack of confidence, we infer a high probability that the model has generated hallucinations. In cases where the model is identified to hallucinate, we will utilize the hidden states from intermediate layers to calibrate the final outputs layers, which is inspired by Chuang et al. (2023). Note, unlike traditional contrastive decoding strategies Li et al. (2023a); Leng et al. (2023); Huang et al. (2023), we do not calibrate all cases. Instead, we focus on mitigating potential hallucination cases to avoid altering non-hallucinatory cases into hallucination ones. Formula 5 show our calibration process.
@@
Note that $\lambda$ is the hyperparameter to control the degree of the intermediate layer. The algorithmic flow can be seen from the pseudo-code below:
@@
As illustrated in Table 5, we conducted experiments using LLaVA-13B. To demonstrate the robustness of our results, we employed two additional relation hallucination datasets, MMRel Nie et al. (2024) and R-bench Wu et al. (2024). During experiments, we set $\lambda$ = 2 ,$\alpha$ = 0.1, $\gamma=0.9$. For a fair comparison, in addition to reporting the baseline model, we also report some training-free methods such as VCD Leng et al. (2023), DoLa Chuang et al. (2023), and OPERA Huang et al. (2023). Our approach achieved improvements across all three relation hallucination datasets. Specifically, on the MMRel dataset, our model achieved a 19.7% improvement compared to the baseline setting.
@@
+ VCD Leng et al. (2023) 38.32 41.96 22.05
+ DoLa Chuang et al. (2023) 36.96 39.68 23.52
+ OPERA Huang et al. (2023) 35.73 39.22 26.73
@@
Table 5: The hallucination rates of LLaVA-13B and its variants(lower is better) with hallucination mitigation methods across Reefknot and two other relation hallucination benchmarks (i.e., MMRel & R-bench).
@@ 6 Case study
To intuitively demonstrate the generation process of hallucination, we visualized a real case from Reefknot in Figure 8. In the image, when we ask “Is the boy eating pizza in the photo?", MLLMs are unable to provide precise answers. To investigate this, we analyze the probabilities associated with the top five predicted tokens at each layer of the language model. As illustrated in Figure 8, the model typically converges on the correct answer as it progresses from shallow to deep layers. Notably, hallucinations didn’t occur until the final layer. Additionally, we can observe that as the number of layers increases, the model increasingly focuses on narrowing down the range of answer choices, with the probability for both ‘yes’ and ‘no’ rising in deeper layers.
However, as depicted in the changes of colors, when it reaches the final decoder block (layer 38-40th), the model’s choices suddenly become uncertain and the entropy rises, accompanied by the emergence of hallucinations. So it is instinctive to utilize the entropy to detect hallucination and use logits of the intermediate layers, which are less likely to hallucinate, to calibrate the final logits to correct the answer. Following our calibration utilizing Equation 5, we not only dispel potential relation-level hallucination but also significantly bolster the confidence of model’s responses.
@@ 7 Related Work
Large Language Models (LLMs) are powerful tools, and exploring their interpretability Han et al. (2024); Huo et al. (2024b), operational mechanisms, and methods to ensure trustworthy responsesHu et al. (2023) has become a crucial area of research. In the realm of MLLMs, hallucinations are typically categorized into three distinct types: object, attribute, and relation, as outlined in prior studies Yin et al. (2024); Liu et al. (2024); Guan et al. (2024). At the object and attribute levels, a considerable number of representative benchmarks such as POPE Li et al. (2023c) and HaELM Wang et al. (2023b) have been introduced by researchers. Many benchmarks are accompanied by a diverse array of evaluation criteria, with generative criteria including CHAIR Li et al. (2023b), THRONE Kaul et al. (2024) and various discriminative criteria being particularly notable. The existing relation hallucinations benchmarks Wu et al. (2024); Nie et al. (2024) focus only on discriminative criteria. The gap of current benchmarks is to identify relationship pairs that accurately reflect the dynamic interactions.
@@
The concept of scene graphs was first introduced by Zhu et al. (2022), offering a novel method to describe and retrieve complex scenes within images. Building on this foundation, Krishna et al. (2016) further advanced the field by introducing the Visual Genome (VG) dataset in 2017. In recent years, there has been significant progress in leveraging this triplet-based structure for various applications. For instance, Yao et al. (2018) utilized scene graphs to generate more detailed image descriptions. Similarly, Li et al. (2019) enhanced VQA systems by incorporating scene graphs, improving the system’s ability to respond to complex questions about images.
@@
Confidence estimation Zou et al. (2023); Chen et al. (2024a) and calibration are essential for enhancing the reliability of LLMs such as GPT-3 Brown et al. (2020). To assess the confidence associated with outputs from LLMs, Kuhn et al. (2023) have developed a method called semantic entropy that utilizes linguistic invariances reflecting shared meanings. However, this method relies on accessing token-level probabilities, which are often unavailable through current black-box APIs. Kadavath et al. (2022) have designed prompts that encourage the models to self-assess their responses and to explicitly calculate the probability that an answer is true; while Lin et al. (2022) have prompted LLMs to provide both an answer and an accompanying confidence level. Furthermore, Manakul et al. (2023) proposed a sampling-based approach for detecting hallucinated facts in the outputs of models.
8 Conclusion
In conclusion, we propose a comprehensive benchmark called Reefknot to evaluate and mitigate relation hallucinations in MLLMs. We construct the dataset with over 20k data through a scene graph based construction pipeline, covering two discriminative tasks (Y/N and MCQ) and one generative task (VQA). Our in-depth evaluation highlights a substantial performance gap on relation hallucination in existing MLLMs, emphasizing the need for more sophisticated reasoning capabilities. Subsequently, we discover that relation hallucinations tend to occur when MLLMs respond with low confidence. Therefore, we propose a Detect-then-Calibrate method to mitigate the relation hallucination via entropy threshold, with an average reduction of 9.75% in the hallucination rate across Reefknot and two other representative relation hallucination datasets. In general, we anticipate that the Reefknot benchmark we propose will serve as a critical foundation for future developments in the field of trustworthy multimodal intelligence, enabling more reliable and robust systems.
9 Limitation
@@
Appendix A Appendix
A.1 Detailed Case Study
@@
A.2 Manual Checking Rules
@@
A.3 More Error Analysis
@@
A.4 Prompt to Categorize
In Figure 14, we show our prompt used to categorize the relation word used in our benchmark during phase(d) in the pipeline of Reefknot.
Prompt: ————————————————– You are a relation term classification assistant. Please help me determine whether the following relational terms/phrases belong to perceptive relationships or cognitive relationships. Perceptive relationships are defined as those involving locational or state-based prepositions, such as “on" and “behind". Cognitive relationships are defined as some words that indicate an action or behavior, such as holding, watching, etc. Here I will give you some demonstrations for reference: {In-context examples} Please answer with a single word in <Perception, Cognition >. Input: {Input}
Figure 14: Prompt template for relation categorization.
A.5 Potential concern
Why detect-than-calibrate method is specific to relation hallucination?
Our approach is specific to relation hallucinations, which uniquely correlate with high entropy, as shown in Table 4 and Figure 7. In contrast, other types of hallucinations do not display this trait. The Table 6 below shows results from the object-level hallucination benchmark POPE Li et al. (2023c), where the average probability is over 90%, resulting in an average entropy of about 0.4, much lower than the $\lambda=0.9$ set in the paper.
Model/Dataset POPE Entropy
LLaVA 0.932 0.358
MiniGPT4-v2 0.904 0.456
Table 6: The average probability and entropy of object-level hallucination instances.
What is the difference between detect-than-calibrate and contrastive decoding based methods like (i.e., vcd , dola)?
Compared to the intuitive contrastive decoding methods, our approach relies on analyzing relational hallucinations and the variation in hidden layers confidence when relation hallucinations occur. In contrast, the intuitive contrastive decoding method does not conduct a detailed analysis of the issues in the individual cases but instead applies a contrastive-decoding approach to all examples, regardless of whether they are hallucination cases. This may mitigate hallucination effects, while leading to non-hallucination examples becoming hallucination ones, which does not fully address the hallucination. In contrast, our method is based on a cause analysis of relation hallucinations from the perspective of model confidence. We first identified a correlation between relation-level hallucinations and high entropy values. Then, we filter hallucination cases based on entropy values before calibration operation, thereby avoiding the potential negative impacts of contrastive decoding methods on normal instances.
In a word, our approach differs from contrastive decoding methods in that we selectively correct potential hallucination examples based on the characteristics of relation hallucinations. The application of the sole contrastive methods tend to convert many non-hallucination cases into hallucination cases, resulting in no significant effects, as demonstrated in Table 5.
A.6 VQA Criterion
Model Perception↓ Cognition↓
DeBERTa GPT-assisted DeBERTa GPT-assisted
LLaVA 0.68 0.93 0.67 0.94
Deepseek 0.70 0.89 0.69 0.81
MiniGPT4 0.61 0.78 0.61 0.83
MinCPM 0.76 0.87 0.69 0.94
Table 7: Comparison of Hallucination Rates between DeBERTa-Based and GPT-Assisted Methods in VQA Settings
Here are the evaluation criteria for VQA questions. We use the DeBERTa [^1] https://huggingface.co/microsoft/deberta-v2-xlarge-mnli model to determine whether the models entail each other. Only when the label and response contain each other will it be judged as a correct reply. We show our key function code in Figure 15. To show potential bias associated with the form of constrained prompt and DeBERTa, we compared the evaluation results using a GPT-assisted method. In this setting, no restrictions were imposed on the prompt (e.g., "What is the relationship between A and B?"). The results, as shown in Table 7, demonstrate that when the responses are not constrained by the prompt, the performance of the vision language models in VQA about relationships is significantly impaired. To further validate the effectiveness of the DeBERTa model, we compared the judgments made by the DeBERTa with those of human checkers, which demonstrate that DeBERTa can accurately assess the correctness of the model’s responses in Table 8. Besides, we show a VQA case in Figure 17.
Ground truth Response DeBERTa Manual check
sunlight is shining on train sunlight is illuminating train 1 1
bear is reading book bear is sitting on book 0 0
picture is hanging wall picture is on wall 1 1
man is lying on couch man is lying couch 1 1
Table 8: The comparison of results between DeBERTa and manual Check for real cases
A.7 More Cases
A.8 Visualization of Relation Word
@@
3 #label is groundtruth;responese is the respone of VLM;